OpenStudy (dmezzullo):

1.05] Use the distributive property to simplify the expression below: -6(2u - 1) -12u + 6 -12u - 6 -12u - 1 12u - 6

OpenStudy (anonymous):

A -6(2u-1) = -6(2u) +6 = -12u + 6

OpenStudy (andriod09):

\[-6(2u-1)\]\[-12u+6\]